Bill can row 3 mph in still water. It takes him 3 hr 36 min to go 3 mi upstream and then return. Find the speed of the current.
Answer by mananth(16946)

Rowing speed = 3 mph
current speed = x mph

against current = 3 - x
with current = 3 + x

Total distance =3 miles Total Time= 3.60 hours
Time upstream + time downstream =

3/(3-x)+3/(3+x )=3.6
3(3+x)+ 3(3-x)= 3.60(9- 1X^2)
9+3x+ 9-3x=32.4-3.6 X^2
18 - -32.4 = -3.6 X^2
-14.4 =-3.6 x^2
-3.6 x^2= -14.4
x^2 = 4
X = +/- 2
Ignore negative value
current speed = 2mph